# Exploring How Audio Effects Alter Emotion with Foundation Models

### v2

@@ 1 Introduction
The relationship between sound and human emotion has been extensively studied across multiple domains, particularly within affective computing and music cognition. Prior research has demonstrated how low-level audio features—such as timbre, tempo, pitch, and rhythm—correlate with emotional responses in listeners [1, 2, 3, 4, 5]. Parallel investigations have examined how environmental auditory conditions, including soundscape complexity and the number of simultaneous sound sources, influence affective perception [6]. However, there remains a notable gap in understanding how audio FX—such as reverberation, distortion, modulation, and dynamic range processing—systematically alter emotional responses in music listening contexts.
Unlike fundamental audio features, audio FX are intentional sound design tools commonly applied in music production and performance to modulate aesthetic and emotional impact. Yet, their role in shaping affective perception has been underexplored. For instance, [7] showed that longer reverberation times in recorded music increase ratings of “sublimity,” including emotional qualities like awe and nostalgia. Similarly, distortion has been linked to heightened arousal and affective intensity, potentially evoking excitement or even aggression [2]. Higher rates of vibrato are linked to elevated arousal levels [8, 9]. Additionally, an increase in the number of perceived sound sources correlates with increased arousal but may reduce pleasantness [6].
@@
At the same time, advances in foundation models—large-scale neural architectures pre-trained on multimodal datasets, have demonstrated an ability to capture high-level semantic and perceptual relationships across domains, including with audio FX [10, 11]. In the music domain, foundation models such as MusicLM [12], MERT [13], Qwen [14], and CLAP [15] have shown promising abilities in generating emotionally expressive audio, learning disentangled latent representations, and enabling cross-modal retrieval tasks [16]. These models implicitly encode rich associations between musical structure, timbre, and affective meaning, making them ideal tools to probe novel hypotheses about music-emotion relationships.
@@ 2 Material
We utilized three datasets capturing both categorical and dimensional emotional annotations: EMOPIA [17], DEAM [18], and witheFlow [5]. The EMOPIA dataset is a multimodal (audio and MIDI) resource focused on perceived emotion in pop piano music, containing 1,087 clips from 387 songs. Each clip is annotated with one of four categorical emotions—Excitement, Anger, Sadness, or Calmness—supporting both clip-level classification and song-level emotion analysis. The DEAM dataset comprises 1,802 excerpts and full songs annotated with valence and arousal as continuous, time-varying values per second, as well as aggregated ratings, along with metadata on duration, genre, and folksonomy tags, enabling regression-based modeling of emotion dynamics. The part of the witheFlow dataset we used includes 235 solo-instrument recordings, annotated by 20 listeners with continuous valence-arousal values per second following the Circumplex Model of Affect [19], as well as multi-label categorical tags based on the Geneva Emotional Music Scale (GEMS-9) [20], making it suitable for both dimensional and categorical emotion modeling in music.
@@
To investigate and interpret musical emotion, we employed three state-of-the-art foundation models, each offering unique capabilities. First, MERT-v1-330M [13] is a 330M-parameter model (24 layers with 1024 hidden units) trained on 160,000 hours of music audio using masked language modeling with RVQ-VAE and CQT teacher signals. MERT operates on 24 kHz audio at 75 frames per second and is particularly suited to capturing fine-grained musical and acoustic emotion cues. Second, CLAP [21] is a dual-encoder model with roughly 630M parameters, combining a Transformer-based (HTS-AT) audio encoder with a RoBERTa-based text encoder. CLAP was trained on 128K–630K audio–text pairs via contrastive learning, using log-Mel spectrogram inputs at 48 kHz to learn a shared audio–text representation space with strong zero-shot generalization. Finally, Qwen2-Audio-7B [22] is a 7B-parameter model featuring a Whisper-style audio encoder and Qwen-style decoder. Qwen2-Audio-7B was trained on over 300K hours of speech, music, and environmental audio to support broad audio-language tasks, including automatic speech recognition, classification, question answering, and conversational interaction.
For the head classifiers of our foundation models, we employed shallow interpretable models tailored to each task and dataset, training them while keeping the foundation model frozen. For regression on DEAM and witheFlow valence-arousal annotations, we used an XGBRegressor [23] to predict continuous emotional dimensions, evaluating performance with MAE, MSE, and $R^{2}$. For single-label classification on EMOPIA, we applied an XGBClassifier, assessing weighted accuracy, precision, recall, and F1-score. For multi-label classification on witheFlow categorical tags, we implemented a OneVsRest strategy with XGBClassifier, measuring F1-micro and F1-macro scores with a threshold of 0.5. This configuration allowed us to effectively adapt gradient-boosted tree models to both dimensional and categorical emotion prediction tasks.
@@ 3.1 Impact on Model Performance
Inspired by [10], that studies the impact of audio effects on foundation models, and [24], that examines performance changes in speech emotion recognition models induced by codec processing, we aimed to explore how incorporating emotional information affects performance in emotion recognition tasks. To investigate this, we trained interpretable machine learning classifiers using embeddings extracted from foundation models across all datasets. This approach allowed us to develop models for multi-class classification, multi-label classification, and regression tasks.
@@ 3.2 Shifts in Emotion Predictions
In Fig. 1, we examine how emotional predictions in the EMOPIA dataset are affected by different audio effects. Notably, high levels of distortion consistently increase Anger predictions while decreasing Calmness across all models [25]. In contrast, delay and chorus effects introduce greater variability, suggesting that these manipulations create ambiguity in the models’ emotional interpretations [26]. Interestingly, increasing the chorus effect tends to boost Calmness predictions in the CLAP and MERT models, whereas increasing the delay effect raises Anger predictions in the CLAP and Qwen models.
@@ 3.3 Embedding Space Alterations
In this section, we present an experiment investigating how the embeddings of a foundation model—representations that encode emotional information—are altered as common audio effects are gradually applied. Inspired by the visualization approach of Deng et al. [10], we identified the most important features from our shallow emotion classification model. Using the top 25 features, we performed UMAP [27] dimensionality reduction and visualized the foundation model’s embeddings for inputs modified at different effect intensities. This approach allowed us to observe how audio effects reshape the structure of the embeddings and to identify which models are most sensitive to these changes.
@@ 4 Conclusions and Future Work
Our study demonstrates that audio effects substantially alter estimated emotion in music. Distortion and phaser, in particular, strongly increase Anger while reducing Calmness, whereas chorus and delay introduce more variability in predictions. Analysis of embedding-space trajectories shows that the magnitude and structure of latent-space shifts reflect the emotional impact of FX, with real-world effect chains producing the most pronounced and coherent changes, likely because they are intentionally designed by artists to evoke strong emotional responses. Among the models, MERT exhibited relative robustness, while CLAP and Qwen were more sensitive to manipulations.

### v3

@@ 1 Introduction
The relationship between sound and human emotion has been extensively studied across multiple domains, particularly within affective computing and music cognition. Prior research has demonstrated how low-level audio features—such as timbre, tempo, pitch, and rhythm—correlate with emotional responses in listeners [laurier2009exploring, panda2020audio, lyberatos2024perceptual, 10944805, lyberatos2025musicinterpretationemotionperception]. Parallel investigations have examined how environmental auditory conditions, including soundscape complexity and the number of simultaneous sound sources, influence affective perception [han2023effect]. However, there remains a notable gap in understanding how audio FX—such as reverberation, distortion, modulation, and dynamic range processing—systematically alter emotional responses in music listening contexts.
Unlike fundamental audio features, audio FX are intentional sound design tools commonly applied in music production and performance to modulate aesthetic and emotional impact. Yet, their role in shaping affective perception has been underexplored. For instance, [wilkie2025reverberation] showed that longer reverberation times in recorded music increase ratings of “sublimity,” including emotional qualities like awe and nostalgia. Similarly, distortion has been linked to heightened arousal and affective intensity, potentially evoking excitement or even aggression [panda2020audio]. Higher rates of vibrato are linked to elevated arousal levels [scherer2015comparing, song2024emotional]. Additionally, an increase in the number of perceived sound sources correlates with increased arousal but may reduce pleasantness [han2023effect].
@@
At the same time, advances in foundation models—large-scale neural architectures pre-trained on multimodal datasets, have demonstrated an ability to capture high-level semantic and perceptual relationships across domains, including with audio FX [deng2025investigating, shah2021all]. In the music domain, foundation models such as MusicLM [agostinelli2023musiclm], MERT [li2023mert], Qwen [qwen2023], and CLAP [elizalde2022clap] have shown promising abilities in generating emotionally expressive audio, learning disentangled latent representations, and enabling cross-modal retrieval tasks [akman2025audio]. These models implicitly encode rich associations between musical structure, timbre, and affective meaning, making them ideal tools to probe novel hypotheses about music-emotion relationships.
@@ 2 Material
We utilized three datasets capturing both categorical and dimensional emotional annotations: EMOPIA [hung2021emopia], DEAM [alajanki2016benchmarking], and witheFlow [lyberatos2025musicinterpretationemotionperception]. The EMOPIA dataset is a multimodal (audio and MIDI) resource focused on perceived emotion in pop piano music, containing 1,087 clips from 387 songs. Each clip is annotated with one of four categorical emotions—Excitement, Anger, Sadness, or Calmness—supporting both clip-level classification and song-level emotion analysis. The DEAM dataset comprises 1,802 excerpts and full songs annotated with valence and arousal as continuous, time-varying values per second, as well as aggregated ratings, along with metadata on duration, genre, and folksonomy tags, enabling regression-based modeling of emotion dynamics. The part of the witheFlow dataset we used includes 235 solo-instrument recordings, annotated by 20 listeners with continuous valence-arousal values per second following the Circumplex Model of Affect [russell1980circumplex], as well as multi-label categorical tags based on the Geneva Emotional Music Scale (GEMS-9) [jacobsen2024assessing], making it suitable for both dimensional and categorical emotion modeling in music.
@@
To investigate and interpret musical emotion, we employed three state-of-the-art foundation models, each offering unique capabilities. First, MERT-v1-330M [li2023mert] is a 330M-parameter model (24 layers with 1024 hidden units) trained on 160,000 hours of music audio using masked language modeling with RVQ-VAE and CQT teacher signals. MERT operates on 24 kHz audio at 75 frames per second and is particularly suited to capturing fine-grained musical and acoustic emotion cues. Second, CLAP [wu2023large] is a dual-encoder model with roughly 630M parameters, combining a Transformer-based (HTS-AT) audio encoder with a RoBERTa-based text encoder. CLAP was trained on 128K–630K audio–text pairs via contrastive learning, using log-Mel spectrogram inputs at 48 kHz to learn a shared audio–text representation space with strong zero-shot generalization. Finally, Qwen2-Audio-7B [chu2024qwen2] is a 7B-parameter model featuring a Whisper-style audio encoder and Qwen-style decoder. Qwen2-Audio-7B was trained on over 300K hours of speech, music, and environmental audio to support broad audio-language tasks, including automatic speech recognition, classification, question answering, and conversational interaction.
For the head classifiers of our foundation models, we employed shallow interpretable models tailored to each task and dataset, training them while keeping the foundation model frozen. For regression on DEAM and witheFlow valence-arousal annotations, we used an XGBRegressor [chen2016xgboost] to predict continuous emotional dimensions, evaluating performance with MAE, MSE, and $R^{2}$. For single-label classification on EMOPIA, we applied an XGBClassifier, assessing weighted accuracy, precision, recall, and F1-score. For multi-label classification on witheFlow categorical tags, we implemented a OneVsRest strategy with XGBClassifier, measuring F1-micro and F1-macro scores with a threshold of 0.5. This configuration allowed us to effectively adapt gradient-boosted tree models to both dimensional and categorical emotion prediction tasks.
@@ 3.1 Impact on Model Performance
Inspired by [deng2025investigating], that studies the impact of audio effects on foundation models, and [albin2015objective], that examines performance changes in speech emotion recognition models induced by codec processing, we aimed to explore how incorporating emotional information affects performance in emotion recognition tasks. To investigate this, we trained interpretable machine learning classifiers using embeddings extracted from foundation models across all datasets. This approach allowed us to develop models for multi-class classification, multi-label classification, and regression tasks.
@@ 3.2 Shifts in Emotion Predictions
In Fig. 1, we examine how emotional predictions in the EMOPIA dataset are affected by different audio effects. Notably, high levels of distortion consistently increase Anger predictions while decreasing Calmness across all models [gao2022black]. In contrast, delay and chorus effects introduce greater variability, suggesting that these manipulations create ambiguity in the models’ emotional interpretations [cesarini2024reverb]. Interestingly, increasing the chorus effect tends to boost Calmness predictions in the CLAP and MERT models, whereas increasing the delay effect raises Anger predictions in the CLAP and Qwen models.
@@ 3.3 Embedding Space Alterations
In this section, we present an experiment investigating how the embeddings of a foundation model—representations that encode emotional information—are altered as common audio effects are gradually applied. Inspired by the visualization approach of Deng et al. [deng2025investigating], we identified the most important features from our shallow emotion classification model. Using the top 25 features, we performed UMAP [mcinnes2018umap] dimensionality reduction and visualized the foundation model’s embeddings for inputs modified at different effect intensities. This approach allowed us to observe how audio effects reshape the structure of the embeddings and to identify which models are most sensitive to these changes.
@@ 4 Conclusions and Future Work
Our study demonstrates that audio effects substantially alter estimated emotion in music. Distortion, in particular, strongly increase Anger while reducing Calmness, whereas chorus and delay introduce more variability in predictions. Analysis of embedding-space trajectories shows that the magnitude and structure of latent-space shifts reflect the emotional impact of FX, with real-world effect chains producing the most pronounced and coherent changes, likely because they are intentionally designed by artists to evoke strong emotional responses. Among the models, MERT exhibited relative robustness, while CLAP and Qwen were more sensitive to manipulations.
@@
5 Acknowledgments
This work was supported by AWS resources, which were provided by the National Infrastructures for Research and Technology GRNET and funded by the EU Recovery and Resiliency Facility